Hello! I need to know how to set up the following question? Thanks so much!
How many ways can a male and female be selected to decorate for a party from a club consisting of 7 members, 5 being men and 2 are women.
There are ______ number of ways that male and female can be selected?

Answer by Maths68(1474)

We need to select 1 man from 5 and 1 woman from 2
Use
nCr=n!/r!(n-r)!
1 man from 5
5C1=5!/1!(5-1)!
5C1=5!/1!4!
5C1=5*4!/1!4!
Cancel out 4!
5C1=5 ways
1 woman from 2
2C1=2!/1!(2-1)!
2C1=2!/1!1!
2C1=2*1
2C1=2 ways
Total Number of ways = 5C2*2C1=5*2=10 ways

One man can be selected out of 5 men = 5 ways
One woman can be select out of 2 women = 2 ways
There are 10 number of ways that male and female can be selected